-7x+y=-19
-2x+3y=-19

Answer by tommyt3rd(5050)

m3: -7x+y=-19
-2x+3y=-19
-21x+3y=-57
-2x+3y=-19

subtracting the equations...

-2x+21x+3y-3y=-19+57
19x=38
x=2
which leads us to y=5

:)